# Integrated population model reveals human and environment driven changes in Baltic ringed seal (Pusa hispida botnica) demography and behavior

### v1

@@ 2.3.2 Survival and mortality
In each year, the mortality sub-process was described as a multi-state transition model, where seals are assigned to one of five states: those that survived ($S$), became hunted in Finland ($H^{\text{fi}}$), became hunted in Sweden either during the spring ($H^{\text{sw(1)}}$) or during the fall ($H^{\text{sw(2)}}$), or died due to other causes ($D$). Due to a lack of data on the magnitude of bycatch, it was included with other sources of mortality.
@@
where $\mu_{f,a}$ is the mortality rate of female seals of age $a$, and $c$ is a parameter that determines how quickly the mortality rate approaches that of adults as the seals age. Male survival rates were modeled in terms of deviations from female survival rates, so that a priori mortality rates were expected to be identical between the sexes, while allowing the data to inform us on potential differences (Appendix S2). The probability, $\phi_{s,a}$, that a seal survives all sources of mortality other than hunting throughout the year is then given by the exponential function
@@
To be able to predict future population responses to alternative management decisions, we explicitly modeled the within-year dynamics of hunting as a function of hunting quotas. We assumed all Finnish hunting takes place in April and May when ringed seals are basking on sea ice, and that the rate at which seals are hunted is proportional to the density of seals hauled-out on ice. We estimated the density of seals along the Finnish coast, in each age and sex group, as $\omega_{s,a,t}n_{s,a,t}/C_{t}$, where $\omega_{s,a,t}$ denotes the proportion of seals expected to be hauled-out on sea ice in year $t$ (see Section 2.3.3), and $C_{t}$ is the ice extent across the Bothnian Bay during peak molting. We further assumed that the number of active hunters in Finland is proportional to the number of unused hunting licenses. The within-year dynamics of hunting in Finland can therefore be expressed as a system of two ordinary differential equations,
@@
where $H_{s,a,t}^{\text{fi}}$ denotes the expected number of seals with sex $s$ and age $a$ that have been hunted in Finland, $Q^{\text{fi}}_{t}$ is the hunting quota for year $t$, $H_{t}^{\text{fi}}(\tau)=\sum_{s,a}H_{s,a,t}^{\text{fi}}(\tau)$ is the total size of the Finnish hunting bag at a time $\tau$ after the onset of the hunting season, $\tilde{n}_{s,a,t}(\tau)$ is the number of seals that are alive at time $\tau$ and $E^{\text{fi}}_{s,a,t}$ is the per capita hunting effort directed towards each demographic group in year $t$.
We assumed that there was stochastic variation in per capita hunting effort across the years, without any temporal trends:
@@
Here, $\hat{E}_{s,a}$ is the median per capita hunting effort, and $\epsilon_{t}\sim\text{N}(0,\sigma_{E_{\text{fi}}}^{2})$ is a noise term. Setting the initial size of the hunting bag to zero, an approximate closed form solution to Equations (6) and (7) can be obtained to estimate the expected composition of the hunting bag, and hence the probability, $\rho^{H^{\text{fi}}}_{s,a,t}$, that a seal with sex $s$ and age $a$ ends up being hunted in Finland during year $t$ (Appendix S3). Swedish hunting during both the spring and the fall were modeled similarly to Finnish hunting, except we did not include any dependence on the extent of sea ice (Appendix S3). Spring hunting in Sweden was assumed to take place during May and June, and fall hunting during September and October (Figure 2).
@@
Note that while the expected size of the hunting bag implied by Equations (6) and (7) depends on the hunting quota, the stochastic model given by Equation (1) does not. This both considerably simplifies the model, and accounts for the possibility of illegal or unrecovered hunting.
2.3.3 Haul-out probability
The proportion of seals that are visible on ice during the molting period depends on the frequency with which seals move into and out of the water. We assumed that during the molting period, seals with sex $s$ and age $a$ move onto the ice at a per-capita rate of $\zeta^{1}_{s,a,t}$, and move back into the water at a per-capita rate of $\zeta^{0}_{s,a,t}$. The proportion of seals hauled-out on ice is then given by
@@ 2.3.3 Haul-out probability
To account for the possibility that a larger fraction of seals haul-out during low ice cover, we modeled the per-capita rate at which hauled-out seals move into the water as a logistic function that decreases as ice cover approaches zero. In other words, we assumed than on average, seals remain on the ice for longer periods of time when sea ice is scarce:
@@
Here, $\hat{\omega}_{s,a}$ is the expected proportion of seals hauled-out when sea ice is abundant, $\alpha_{0}$ and $\alpha_{1}$ are parameters that determine the shape of the logistic curve, and $d$ is the ratio between the lower and upper asymptotes of $\zeta^{0}_{s,a,t}$. A non-zero lower asymptote for $\zeta^{0}_{s,a,t}$ accounts for the fact that the maximum amount of time seals can spend on ice may be constrained by the need to forage or cool off.
We assumed that on average sub-adult and adult seals haul-out with the same probability (i.e. $\hat{\omega}_{s,a}=\hat{\omega},\ \forall a>0$), but note that during the time of aerial surveys sub-adults are typically thought to be underrepresented on ice compared to adults (smith1970). Hence, $\hat{\omega}$ should be regarded as the average haul-out probability across sub-adults and adults. Pups largely complete molting by the time aerial surveys are conducted, and are significantly less likely to be on ice (smith1970; lydersen1993). Thus, we assumed that the typical haul-out probability of pups is a constant fraction $\delta\in(0,1)$ of the haul-out probability of sub-adults and adults (i.e. $\hat{\omega}_{s,0}=\delta\hat{\omega})$.
@@ 2.3.4 Reproduction
To account for the diminishing effects of organochlorine contamination on the reproductive rates of ringed seals, we modeled the probability that an adult female gives birth in the absence of density dependent effects as a time-varying logistic function increasing from a historic low of $b_{\text{min}}$ to a theoretical maximum of $b_{\text{max}}$, so that
@@
where $N_{t-1}=\sum_{s,a}n_{s,a,t-1}$ is the total population size during the previous year, $\theta_{0}$ is the average failure rate of pregnancies at zero population density and $\theta_{1}$ is the rate at which such failed pregnancies increase with population size (Appendix S4).
@@ 2.4.1 Aerial survey estimates
Here, $\mathbf{\tilde{n}}_{t}$ is the expected population state at the time of aerial surveys (Section 2.3.2), $\boldsymbol{\omega}_{t}$ is a vector of haul-out probabilities (Section 2.3.3) and $r$ is a variance parameter. The negative binomial distribution is commonly used to model biological count data, and can be thought of as a Poisson distribution with a rate parameter that varies stochastically as a result of, e.g., changes in observability or demographic stochasticity on $\mathbf{\tilde{n}}_{t}$ that lead to over-dispersion in the counts (Linden2011).
2.4.2 Hunting bag and hunting samples
The total size of the hunting bag for both Finland and Sweden are known exactly in principle, but we nonetheless assumed a small ($\approx 5\%$ CV) variation around them because, in addition to having computational benefits, it accounts for the possibility of unrecovered or illegally hunted seals, as well as rare instances of misreporting:
$y_{t}^{\text{hb}}|\mathbf{u}^{(1)}_{t}\sim N(H_{t},(0.05H_{t})^{2}).$ (17)
Here, $y_{t}^{\text{hb}}$ denotes the observed size of the hunting bag and $H_{t}$ denotes the total number of seals hunted in either Finland or Sweden during year $t$. We used separate observation models for spring and fall hunting in Sweden, except in 2015 and 2016 when only the total size of the hunting bag was known.
Assuming seals are sampled independently from the hunting bag, the age and sex composition of the sampled seals from Finland and Sweden can be modeled using the multinomial distribution
@@ 2.4.2 Hunting bag and hunting samples
where $\mathbf{y}_{t}^{\text{hs}}$ is the demographic composition of the hunting samples in year $t$, $\mathbf{u}^{(1)}_{H,t}$ is the composition of the hunting bag, $v_{s,a}$ is the relative probability that a seal with sex $s$ and age $a$ is sampled from the hunting bag, and $\odot$ denotes the element wise, or Hadamard product. We assumed that seals are sampled randomly from the hunting bag in Finland, and in Sweden during the fall (i.e. $v_{s,a}=1,\ \forall s,a$). However, sampling during the spring has not been random in Sweden since larger seals are specifically requested to assess reproductive health. The observation model to estimate $\mathbf{v}$ for spring hunting in Sweden is described in Appendix S5.
Observation models for samples that were missing either sex or age information are described in Appendix S6.
@@ 2.4.3 By-catch samples
where $\mathbf{y}_{t}^{\text{bs}}$ is the demographic composition of the bycaught samples in year $t$, and $\boldsymbol{\psi}$ is a 12-simplex of weights accounting for possible deviations between the composition of bycatch and sources of mortality other than hunting (Appendix S2).
Observation models for samples that were missing either sex or age information are presented in Appendix S6.
@@ 2.4.4 Observations of reproductive success
We assumed that the reproductive status of sampled females was representative of the population as a whole. Moreover, we assumed that the failure rate of pregnancies is highest prior to the implantation of embryos, and decreases linearly to zero at parturition. Assuming no errors in the detection of fetuses, the number of observed pregnancies during the fall of year $t$ can be modeled using the binomial distribution
@@
$p_{t}=b_{t+1}e^{(1-\tau_{p}(2-\tau_{p}))\theta_{0}e^{\theta_{1}N_{t}}}$ (21)
is the pregnancy rate of the population, and $\tau_{p}\approx 0.5$ is the average time, in years, between mating and the sampling of pregnant females (Appendix S4).
@@
The probabilities $\gamma^{ij}_{t}$ for each assessment outcome will depend on the birth rate $b_{t}$, the probability $\kappa$ that a seal that has not given birth has a CA, and the detection probabilities for placental scars, $\pi_{s}$, and CA, $\pi_{c}$, respectively. Hence, for a randomly sampled seal both a placental scar and a CA are observed ($z^{11}_{k,t}=1$) with probability $\gamma_{t}^{11}=b_{t}\pi_{s}\pi_{c}$, corresponding to sampling a seal that has recently given birth and successfully detecting both the placental scar and the CA. If a placental scar is observed without a corresponding CA ($z^{10}_{k,t}=1$), it must be that a CA was present but not seen. Such an outcome will occur with probability $\gamma_{t}^{10}=b_{t}\pi_{s}(1-\pi_{c})$. Finally, if a CA is observed without a corresponding placental scar ($z^{01}_{k,t}=1$), it means that either the female recently gave birth but the placental scar was not detected (i.e. because it faded), or the female did not recently give birth but nonetheless had a CA which was detected. Such an assessment will occur with probability $\gamma_{t}^{01}=b_{t}(1-\pi_{s})\pi_{c}+(1-b_{t})\kappa\pi_{c}$ The probability that neither a placental scar nor a CA is observed is then $\gamma_{t}^{00}=1-\gamma_{t}^{11}-\gamma_{t}^{10}-\gamma_{t}^{01}$.
@@
An extension of the observation model described in this section to handle incomplete data is presented in Appendix S6.
@@ 2.5.1 Posterior inference
Following standard practice in IPMs, we assumed the different data sources were independent and constructed the joint likelihood as a product of the individual data likelihoods presented in Section 2.4 (abadi2010). Prior distributions for model parameters are given in Appendix S2. Posterior inference was conducted using Markov chain Monte Carlo (MCMC) methods and all analyses were performed using R Statistical Software version 4.2.0 (Rteam2021). The Markov chain sampling was implemented using Stan version 2.26.23 (stan2018), a probabilistic programming language that uses gradient-based methods to draw samples from the posterior distributions of the parameters. Because Stan relies on gradient information, it does not permit discrete valued parameters. As a result, all discrete valued parameters in our model, along with their corresponding probability distributions, were approximated using continuous numbers and distributions (Appendix S7).
@@
We simulated four Markov Chains, and drew 4,000 posterior samples with each chain. The first 2,000 samples drawn from each chain were treated as the warm-up period and were discarded. The resulting Effective Sample Size (stan2018) was greater than 1,500 for all model parameters. Convergence checks were performed using the R-hat diagnostic (stan2018) as well as visual inspection of trace plots. All R-hat values were less than 1.01, and the trace plots for each parameter appeared to be well-mixed, indicating that the model had converged.
Model fit was assessed using posterior predictive checks (Figure 6a and Appendix S1). A visual comparison of the observed distribution of data with the distribution of simulated data showed no signs of model misfit.
2.5.2 Life Table analysis
Our modeling approach can be viewed as a generalization of a Leslie matrix model that incorporates demographic stochasticity (caswell2001; Buckland2004). The contribution of demographic stochasticity to population dynamics diminishes as population size increases. With thousands of ringed seals in the Bothnian Bay, the dynamics can be approximated using the conditional expectation given by the matrix projection model
$\text{E}[\mathbf{n}_{t+1}|\mathbf{n}_{t}]=\mathbf{L}_{t}(N_{t})\mathbf{n}_{t},$ (24)
where the matrix $\mathbf{L}_{t}$ is often called the Leslie matrix in the case of age-structured populations.
Matrix models provide a convenient tool for life table analysis (caswell2001). For instance, when the matrix $\mathbf{L}$ is constant over time (e.g. when analyzing linearized dynamics far from the carrying capacity), the asymptotic population growth rate is given by its dominant eigenvalue $\lambda$, and the stable demographic structure is given by its dominant right eigenvector (caswell2001). The dominant eigenvalue can be obtained numerically from the Euler-Lotka Equation
.$\frac{b}{2}\frac{\lambda^{-5}\prod_{i=0}^{4}\phi_{f,i}}{1-\phi_{f,5+}/\lambda}=1$ (25)
where $b$ is the birth rate assumed to be a constant (caswell2001). When $\lambda=1$, the left hand side of Equation (25) yields the net reproduction rate of the population, defined as the expected number of female offspring produced by a female seal throughout its lifetime (caswell2001).

### v2

@@ 2.3.2 Survival and mortality
In each year, the mortality sub-process was described as a multi-state transition model, where seals are assigned to one of five states: those that survived ($S$), were harvested in Finland ($H^{\text{fi}}$), were harvested in Sweden either during the spring ($H^{\text{sw(1)}}$) or during the fall ($H^{\text{sw(2)}}$), or died due to other causes ($D$). Due to a lack of data on the magnitude of bycatch, it was included with other sources of mortality.
@@
where $\mu_{f,a}$ is the mortality rate of female seals of age $a$, and $c$ is a parameter that determines how quickly the mortality rate approaches that of adults as the seals age. Male survival rates were modeled in terms of deviations from female survival rates, so that a priori mortality rates were expected to be identical between the sexes, while allowing the data to inform us on potential differences (Supplement 2). The probability, $\phi_{s,a}$, that a seal survives all sources of mortality other than hunting throughout the year is then given by the exponential function
@@
To be able to predict future population responses to alternative management decisions, we explicitly modeled the within-year dynamics of hunting as a function of hunting quotas. We assumed all Finnish hunting takes place in April and May when ringed seals are basking on sea ice, and that the harvest rate is proportional to the density of seals hauled out on ice. We estimated the density of seals along the Finnish coast, in each age and sex group, as $\omega_{s,a,t}n_{s,a,t}/C_{t}$, where $\omega_{s,a,t}$ denotes the proportion of seals expected to be hauled out on sea ice in year $t$ (see Section 2.3.3), and $C_{t}$ is the ice extent across the Bothnian Bay during peak molting. We further assumed that the number of active hunters in Finland is proportional to the number of unused hunting licenses, implying a Holling Type I functional response in the absence of a quota (holling1959). The within-year dynamics of hunting in Finland can therefore be expressed as a system of two ordinary differential equations,
@@
where $H_{s,a,t}^{\text{fi}}$ denotes the expected number of seals with sex $s$ and age $a$ that have been harvested in Finland, $Q^{\text{fi}}_{t}$ is the hunting quota for year $t$, $H_{t}^{\text{fi}}(\tau)=\sum_{s,a}H_{s,a,t}^{\text{fi}}(\tau)$ is the total harvest in Finland at a time $\tau$ after the onset of the hunting season, $\tilde{n}_{s,a,t}(\tau)$ is the number of seals that are alive at time $\tau$ and $E^{\text{fi}}_{s,a,t}$ is the per capita harvest rate directed towards each demographic group in year $t$.
We assumed that there was stochastic variation in per capita harvest rates across the years, without any temporal trends:
@@
Here, $\hat{E}_{s,a}$ is the median per capita harvest rate, and $\epsilon_{t}\sim\text{N}(0,\sigma_{E_{\text{fi}}}^{2})$ is a noise term. Setting the initial harvest to zero, an approximate closed form solution to Equations (6) and (7) can be obtained to estimate the expected composition of the harvests, and hence the probability, $\rho^{H^{\text{fi}}}_{s,a,t}$, that a seal with sex $s$ and age $a$ is harvested in Finland during year $t$ (Supplement 3). Swedish hunting during both the spring and the fall were modeled similarly to Finnish hunting. However, unlike hunting in Finland which takes place almost entirely on sea ice, protective hunting in Sweden is likely to be opportunistic. We therefore assumed that the harvest rate in Sweden was proportional to the total number of ringed seals in the Bothnian Bay, rather than the density of ringed seals on sea ice (Supplement 3). Spring hunting in Sweden was assumed to take place during May and June, and fall hunting during September and October (Figure 2).
@@
Note that while the expected harvests, as implied by Equations (6) and (7), are constrained by the hunting quotas, the realized harvests given by the stochastic model in Equation (1) may exceed these quotas. This formulation not only simplifies the model considerably but also accounts for the possibility of illegal or unrecovered hunting.
2.3.3 Haul out probability
The proportion of seals that are visible on ice during the molting period depends on the frequency with which seals move into and out of the water. We assumed that during the molting period, seals with sex $s$ and age $a$ move onto the ice at a per-capita rate of $\zeta^{1}_{s,a,t}$, and move back into the water at a per-capita rate of $\zeta^{0}_{s,a,t}$. The proportion of seals hauled out on ice is then given by
@@ 2.3.3 Haul out probability
Although our assumption of constant movement rates implies exponentially distributed haul out and foraging durations, Equation (10) remains valid for any distribution with a finite expectation (janssen2006).
@@
To account for the possibility that a larger fraction of seals haul out during low ice cover, we modeled the per-capita rate at which hauled-out seals move into the water as a logistic function that decreases as ice cover approaches zero. In other words, we assumed than on average, seals remain on the ice for longer periods of time when sea ice is scarce:
@@
Here, $\hat{\omega}_{s,a}$ is the expected proportion of seals hauled out when sea ice is abundant, $\alpha_{0}$ and $\alpha_{1}$ are parameters that determine the shape of the logistic curve, and $d$ is the ratio between the lower and upper asymptotes of $\zeta^{0}_{s,a,t}$. A non-zero lower asymptote for $\zeta^{0}_{s,a,t}$ accounts for the fact that the maximum amount of time seals can spend on ice may be constrained by the need to forage or cool off.
We assumed that on average sub-adult and adult seals haul out with the same probability (i.e. $\hat{\omega}_{s,a}=\hat{\omega},\ \forall a>0$), but note that during the time of aerial surveys sub-adults are typically thought to be underrepresented on ice compared to adults (smith1970). Hence, $\hat{\omega}$ should be regarded as the average haul out probability across sub-adults and adults. Pups largely complete molting by the time aerial surveys are conducted, and are significantly less likely to be on ice (smith1970; lydersen1993). Thus, we assumed that the baseline haul out probability of pups is a constant fraction $\delta\in[0,1]$ of that for sub-adults and adults (i.e. $\hat{\omega}_{s,0}=\delta\hat{\omega}$), with the effect of sea ice modeled in the same way as for sub-adults and adults using Equations (10)-(12).
@@ 2.3.4 Reproduction
To account for the diminishing effects of organochlorine contamination on the reproductive rates of ringed seals, we modeled the probability that an adult female gives birth to a single pup in the absence of density dependent effects as a time-varying logistic function increasing from a historic low of $b_{\text{min}}$ to a theoretical maximum of $b_{\text{max}}$, so that
@@
where $N_{t-1}=\sum_{s,a}n_{s,a,t-1}$ is the total population size during the previous year, $\theta_{0}$ is the average failure rate of pregnancies at zero population density and $\theta_{1}$ is the rate at which such failed pregnancies increase with population size (Supplement 4).
@@ 2.4.1 Aerial survey estimates
Here, $\mathbf{\tilde{n}}_{t}$ is the expected population state at the time of aerial surveys (Section 2.3.2), $\boldsymbol{\omega}_{t}$ is a vector of haul out probabilities (Section 2.3.3) and $r$ is a variance parameter. The negative binomial distribution is commonly used to model biological count data, and can be thought of as a Poisson distribution with a rate parameter that varies stochastically as a result of, e.g., changes in observability or demographic stochasticity on $\mathbf{\tilde{n}}_{t}$ that lead to over-dispersion in the counts (Linden2011).
2.4.2 Harvest totals and samples
The total harvests for both Finland and Sweden are known exactly in principle, but we nonetheless assumed a small ($\approx 5\%$ CV) variation around them because, in addition to having computational benefits, it accounts for the possibility of unrecovered or illegally hunted seals, as well as rare instances of misreporting:
$y_{t}^{\text{ht}}|\mathbf{u}^{(1)}_{t}\sim N(H_{t},(0.05H_{t})^{2}).$ (17)
Here, $y_{t}^{\text{ht}}$ denotes the observed harvest totals and $H_{t}$ denotes the actual harvest totals in either Finland or Sweden during year $t$. We used separate observation models for spring and fall hunting in Sweden, except in 2015 and 2016 when only the total harvest was known.
Assuming that sampling of seals from the harvests is independent, the age and sex composition of the sampled seals from Finland and Sweden can be modeled using the multinomial distribution
@@ 2.4.2 Harvest totals and samples
where $\mathbf{y}_{t}^{\text{hs}}$ is the demographic composition of the recovered samples in year $t$, $\mathbf{u}^{(1)}_{H,t}$ is the composition of the harvests, $v_{s,a}$ is the relative probability that a seal with sex $s$ and age $a$ is sampled from the harvests, and $\odot$ denotes the element wise, or Hadamard product. We assumed that seals are sampled randomly from the harvests in Finland, and in Sweden during the fall (i.e. $v_{s,a}=1,\ \forall s,a$). However, sampling during the spring has not been random in Sweden since larger seals are specifically requested to assess reproductive health. The observation model to estimate $\mathbf{v}$ for spring hunting in Sweden is described in Supplement 5. Observation models for samples that were missing either sex or age information are described in Supplement 6.
@@ 2.4.3 By-catch samples
where $\mathbf{y}_{t}^{\text{bs}}$ is the demographic composition of the bycaught samples in year $t$, and $\boldsymbol{\psi}$ is a 12-simplex of weights accounting for possible deviations between the composition of bycatch and sources of mortality other than hunting (Supplement 2).
Observation models for samples that were missing either sex or age information are presented in Supplement 6.
@@ 2.4.4 Observations of reproductive success
Given reports of significant late-term pregnancy losses in other pinniped populations and the increased energetic demands of late-stage gestation (pitcher1998; mckenzie2005; stenson2016), we assumed the rate of pregnancy losses increased linearly from zero at conception to a maximum at parturition. However, we note that there is little evidence of late-term pregnancy losses in Baltic ringed seals, and it has been suggested that the influence of nutritional factors on pinniped reproduction is greatest at the point of implantation (boyd1991). Hence, for sensitivity analysis we also considered a model where the rate of pregnancy losses was assumed to decrease linearly (Supplement 4).
Assuming the reproductive status of sampled females is representative of the population as a whole, and there are no errors in the detection of fetuses, the number of observed pregnancies during the fall of year $t$ can be modeled using the binomial distribution
@@
$p_{t}=b_{t+1}e^{(1-\tau_{p}^{2})\theta_{0}e^{\theta_{1}N_{t}}}$ (21)
is the pregnancy rate of the population, and $\tau_{p}\approx 0.5$ is the average time, in years, between mating and the sampling of pregnant females (Supplement 4).
@@
The probabilities $\gamma^{ij}_{t}$ for each assessment outcome will depend on the birth rate $b_{t}$, the probability $\kappa$ that a seal that has not given birth has a CA, and the detection probabilities for placental scars, $\pi_{s}$, and CA, $\pi_{c}$, respectively. Hence, for a randomly sampled seal both a placental scar and a CA are observed ($z^{11}_{k,t}=1$) with probability $\gamma_{t}^{11}=b_{t}\pi_{s}\pi_{c}$, corresponding to sampling a seal that has recently given birth and successfully detecting both the placental scar and the CA. If a placental scar is observed without a corresponding CA ($z^{10}_{k,t}=1$), it must be that a CA was present but not seen. Such an outcome will occur with probability $\gamma_{t}^{10}=b_{t}\pi_{s}(1-\pi_{c})$. Finally, if a CA is observed without a corresponding placental scar ($z^{01}_{k,t}=1$), it means that either the female recently gave birth but the placental scar was not detected (i.e. because it faded), or the female did not recently give birth but nonetheless had a CA which was detected. Such an assessment will occur with probability $\gamma_{t}^{01}=b_{t}(1-\pi_{s})\pi_{c}+(1-b_{t})\kappa\pi_{c}$. The probability that neither a placental scar nor a CA is observed is then $\gamma_{t}^{00}=1-\gamma_{t}^{11}-\gamma_{t}^{10}-\gamma_{t}^{01}$.
@@
An extension of the observation model described in this section to handle incomplete data is presented in Supplement 6.
@@ 2.5.1 Posterior inference
Following standard practice in IPMs, we assumed the different data sources were independent and constructed the joint likelihood as a product of the individual data likelihoods presented in Section 2.4 (abadi2010). Prior distributions for model parameters are given in Supplement 2. Posterior inference was conducted using Markov chain Monte Carlo (MCMC) methods implemented using Stan version 2.26.23 (stan2018), a probabilistic programming language that uses gradient-based methods to draw samples from the posterior distributions of the parameters. Because Stan relies on gradient information, it does not permit discrete valued parameters. As a result, all discrete valued parameters in our model, along with their corresponding probability distributions, were approximated using continuous numbers and distributions (Supplement 7).
@@
We simulated four Markov Chains, and drew 4,000 posterior samples with each chain. The first 2,000 samples drawn from each chain were treated as the warm-up period and were discarded. The resulting Effective Sample Size (stan2018) was greater than 1,500 for all model parameters. Convergence checks were performed using the R-hat diagnostic (stan2018) as well as visual inspection of trace plots. All R-hat values were less than 1.01, and the trace plots for each parameter appeared to be well-mixed, indicating that the model had converged. Model fit was assessed using posterior predictive checks (Supplement 1). A visual comparison of the observed distribution of data with the distribution of simulated data showed no signs of model misfit.
2.5.2 Sensitivity analyses
To examine how different data sources influenced posterior estimates of individual parameters, we refitted our model using various subsets of the available data (gelman2020). We focused on aerial survey data, reproduction data, harvest totals, and dead samples. For each of the four data types, we refitted the model using only data from odd years. For aerial survey data, we also ran models with only pre- and post-2012 data, corresponding respectively to periods with low and high variability in the observed counts. For reproduction data, we ran an additional model with only post-2007 data to determine if long-term reproductive trends could be estimated without direct observations of reproductive signs. In total, we fit our model to eight different datasets: the full dataset and seven reduced versions. For all models, we created prior-posterior comparison plots of key parameters and calculated prior-posterior overlaps (Supplement 1: Figures S5-S9). For models using the reduced datasets, we also computed overlaps of posteriors with those from the full model (Supplement 1: Table LABEL:table:post_post_overlap).
Additionally, we computed Pareto-k statistics for each data type using Pareto smoothed importance sampling, implemented via the R package ’loo’ (loo), in order to assess their relative influence on overall model fit (vehtari2017). The results of these analyses are presented in Supplement 1.
We also ran prior sensitivity analyses for the baseline haul out probability parameter $\hat{\omega}$, which plays a key role in estimating total population size, and was assigned an informative prior in our model due to a lack of direct data (Section 2.3.3). We refitted our model with two alternative priors, each shifted by $0.1$ to the left and right from the original, and compared the resulting estimates for key model parameters.
All models were fitted and their sampling convergence assessed as described in Section 2.5.1. In four of the sensitivity analyses, one of the chains showed poor mixing and was therefore removed from consideration.
2.5.3 Demographic analysis
Since asymptotic demographic analysis is of interest to practical management of Baltic ringed seals, we applied analytical methods for matrix population models to our demographic parameter estimates to calculate key asymptotic quantities, such as the intrinsic population growth rate, stable stage structure, and critical harvest levels that would lead to population decline (caswell2001; ersalman2024). All demographic analyses were performed in R Statistical Software version 4.2.0 (Rteam2021).